# ChromFound: Towards A Universal Foundation Model for Single-Cell Chromatin Accessibility Data

### v2

@@ I.1.1 Adjusted Rand Index (ARI)
ranging from $-1$ to $1$, where higher values indicate better agreement between clustering results and ground-truth labels.
@@ I.1.2 Fowlkes-Mallows Index (FMI)
where TP, FP, and FN are defined as above. It ranges from 00 to $1$, with higher values indicating better clustering performance.
@@ I.1.3 Normalized Mutual Information (NMI)
where $I(\cdot;\cdot)$ is mutual information and $H(\cdot)$ is entropy. It ranges from 00 to $1$, with higher values indicating greater alignment between the clustering and ground-truth labels.
@@ I.1.4 Adjusted Mutual Information (AMI)
ranging from 00 to $1$, where higher values indicate a more robust agreement between clustering and ground-truth, especially in datasets with many clusters.
@@ I.2.1 Biological Conservation Metrics
where $I(\cdot;\cdot)$ is mutual information and $H(\cdot)$ is entropy. It ranges from 00 to $1$, with higher values indicating better alignment with ground-truth cell types.
@@
where $\text{RI}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ and TP, TN, FP, FN are defined as in the clustering section. It ranges from $-1$ to $1$, with higher values indicating better clustering agreement with cell types.
@@
where $a(i)$ is the average distance of cell $i$ to others in its cell type, and $b(i)$ is the smallest average distance to another cell type. It ranges from $-1$ to $1$, with higher values indicating better preservation of biological clustering.
@@
averaging the biological conservation metrics, ranging from 00 to $1$, with higher values indicating better preservation of biological signals.
@@ I.2.2 Batch Mixing Metrics
where $a(i)$ and $b(i)$ are computed with respect to batch labels, with $a(i)$ as the average distance to other cells in the same batch and $b(i)$ as the smallest average distance to cells in another batch. It ranges from 00 to $1$, with higher values indicating better batch mixing.
@@
where $\text{LCC}(G_{c}^{\text{kNN}})$ is the size of the largest connected component in the k-nearest neighbors (kNN) graph of cells in cell type $c$, and $N_{c}$ is the number of cells in $c$. It ranges from 00 to $1$, with higher values indicating better connectivity across batches.
@@
averaging the batch mixing metrics, ranging from 00 to $1$, with higher values indicating better batch correction.
@@ I.3.1 Accuracy
where $\text{tp}_{c}$ is the number of true positives for cell type $c$, and $N_{c}$ is the total number of cells in $c$. It ranges from 00 to $1$, with higher values indicating better classification performance.
@@ I.3.2 Macro F1 Score
[…]
where $\text{Precision}_{c}=\frac{\text{tp}_{c}}{\text{tp}_{c}+\text{fp}_{c}}$ and $\text{Recall}_{c}=\frac{\text{tp}_{c}}{\text{tp}_{c}+\text{fn}_{c}}$, with $\text{fp}_{c}$ and $\text{fn}_{c}$ as false positives and false negatives for cell type $c$. It ranges from 00 to $1$, with higher values indicating better balanced performance across classes.
@@ I.4.1 Pearson Correlation Coefficient (PCC)
where $\bar{x}$ and $\bar{y}$ are the means of the predicted and true values, respectively. It ranges from $-1$ to $1$, with higher (closer to $1$) values indicating stronger positive linear correlation.
@@ I.4.2 Concordance Correlation Coefficient (CCC)
where $\rho$ is the PCC, $\sigma_{x}$ and $\sigma_{y}$ are the standard deviations of the predicted and true values, and $\mu_{x}$ and $\mu_{y}$ are their means. It ranges from $-1$ to $1$, with higher (closer to $1$) values indicating better agreement in both correlation and scale.
@@ I.5.1 Area Under the Receiver Operating Characteristic Curve (AUC-ROC)
where TP, TN, FP, and FN are true positives, true negatives, false positives, and false negatives, respectively. The Area Under the ROC Curve (AUC-ROC) quantifies the overall performance across all thresholds, ranging from 00 to $1$, with higher values indicating better classification performance, where $1$ represents perfect classification and $0.5$ represents random guessing.

### v3

@@ I.1.1 Adjusted Rand Index (ARI)
ranging from 1-1 to $1$, where higher values indicate better agreement between clustering results and ground-truth labels.
@@ I.1.2 Fowlkes-Mallows Index (FMI)
where TP, FP, and FN are defined as above. It ranges from 0 to $1$, with higher values indicating better clustering performance.
@@ I.1.3 Normalized Mutual Information (NMI)
where $I(\cdot;\cdot)$ is mutual information and $H(\cdot)$ is entropy. It ranges from 0 to $1$, with higher values indicating greater alignment between the clustering and ground-truth labels.
@@ I.1.4 Adjusted Mutual Information (AMI)
ranging from 0 to $1$, where higher values indicate a more robust agreement between clustering and ground-truth, especially in datasets with many clusters.
@@ I.2.1 Biological Conservation Metrics
where $I(\cdot;\cdot)$ is mutual information and $H(\cdot)$ is entropy. It ranges from 0 to $1$, with higher values indicating better alignment with ground-truth cell types.
@@
where $\text{RI}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ and TP, TN, FP, FN are defined as in the clustering section. It ranges from 1-1 to $1$, with higher values indicating better clustering agreement with cell types.
@@
where $a(i)$ is the average distance of cell $i$ to others in its cell type, and $b(i)$ is the smallest average distance to another cell type. It ranges from 1-1 to $1$, with higher values indicating better preservation of biological clustering.
@@
averaging the biological conservation metrics, ranging from 0 to $1$, with higher values indicating better preservation of biological signals.
@@ I.2.2 Batch Mixing Metrics
where $a(i)$ and $b(i)$ are computed with respect to batch labels, with $a(i)$ as the average distance to other cells in the same batch and $b(i)$ as the smallest average distance to cells in another batch. It ranges from 0 to $1$, with higher values indicating better batch mixing.
@@
where $\text{LCC}(G_{c}^{\text{kNN}})$ is the size of the largest connected component in the k-nearest neighbors (kNN) graph of cells in cell type $c$, and $N_{c}$ is the number of cells in $c$. It ranges from 0 to $1$, with higher values indicating better connectivity across batches.
@@
averaging the batch mixing metrics, ranging from 0 to $1$, with higher values indicating better batch correction.
@@ I.3.1 Accuracy
where $\text{tp}_{c}$ is the number of true positives for cell type $c$, and $N_{c}$ is the total number of cells in $c$. It ranges from 0 to $1$, with higher values indicating better classification performance.
@@ I.3.2 Macro F1 Score
$\text{Macro F1}=\frac{1}{|C|}\sum_{c\in C}F1_{c},\quad F1_{c}=\frac{2\cdot\text{Precision}_{c}\cdot\text{Recall}_{c}}{\text{Precision}_{c}+\text{Recall}_{c}},$
where $\text{Precision}_{c}=\frac{\text{tp}_{c}}{\text{tp}_{c}+\text{fp}_{c}}$ and $\text{Recall}_{c}=\frac{\text{tp}_{c}}{\text{tp}_{c}+\text{fn}_{c}}$, with $\text{fp}_{c}$ and $\text{fn}_{c}$ as false positives and false negatives for cell type $c$. It ranges from 0 to $1$, with higher values indicating better balanced performance across classes.
@@ I.4.1 Pearson Correlation Coefficient (PCC)
where $\bar{x}$ and $\bar{y}$ are the means of the predicted and true values, respectively. It ranges from 1-1 to $1$, with higher (closer to $1$) values indicating stronger positive linear correlation.
@@ I.4.2 Concordance Correlation Coefficient (CCC)
where $\rho$ is the PCC, $\sigma_{x}$ and $\sigma_{y}$ are the standard deviations of the predicted and true values, and $\mu_{x}$ and $\mu_{y}$ are their means. It ranges from 1-1 to $1$, with higher (closer to $1$) values indicating better agreement in both correlation and scale.
@@ I.5.1 Area Under the Receiver Operating Characteristic Curve (AUC-ROC)
where TP, TN, FP, and FN are true positives, true negatives, false positives, and false negatives, respectively. The Area Under the ROC Curve (AUC-ROC) quantifies the overall performance across all thresholds, ranging from 0 to $1$, with higher values indicating better classification performance, where $1$ represents perfect classification and $0.5$ represents random guessing.